# LLM-guided Chemical Process Optimization with a Multi-Agent Approach

### v1

@@ Abstract
Chemical process optimization is crucial to maximize production efficiency and economic performance. Traditional methods, including gradient-based solvers, evolutionary algorithms, and parameter grid searches, become impractical when operating constraints are ill-defined or unavailable, requiring engineers to rely on subjective heuristics to estimate feasible parameter ranges. To address this constraint definition bottleneck, we present a multi-agent framework of large language model (LLM) agents that autonomously infer operating constraints from minimal process descriptions, then collaboratively guide optimization using the inferred constraints. Our AutoGen-based agentic framework employs OpenAI’s o3 model, with specialized agents for constraint generation, parameter validation, simulation execution, and optimization guidance. Through two phases - autonomous constraint generation using embedded domain knowledge, followed by iterative multi-agent optimization - the framework eliminates the need for predefined operational bounds. Validated on the hydrodealkylation process across cost, yield, and yield-to-cost ratio metrics, the framework demonstrated competitive performance with conventional optimization methods while achieving better computational efficiency, requiring fewer iterations to converge. Our approach converged in under 20 minutes, achieving a 31-fold speedup over grid search. Beyond computational efficiency, the framework’s reasoning-guided search demonstrates sophisticated process understanding, correctly identifying utility trade-offs, and applying domain-informed heuristics. This approach shows significant potential for optimization scenarios where operational constraints are poorly characterized or unavailable, particularly for emerging processes and retrofit applications.
[…]
{tocentry}[Uncaptioned image]
@@ 1 Introduction
Efficient optimization of chemical processes is a central objective in process systems engineering, with direct implications for production efficiency and economic performance1. Conventional optimization approaches include gradient-based solvers such as Interior Point OPTimizer (IPOPT)2 and Sparse Nonlinear OPTimizer (SNOPT)3, which depend on the smoothness and differentiability of the underlying objective function4. When these requirements cannot be met, alternative methods such as evolutionary algorithms5, and derivative-free methods6 provide viable solutions. Furthermore, exhaustive search techniques7 such as grid search systematically evaluate all combinations of discretized parameter values throughout the design space.
@@
However, these optimization methods require a complete and well-defined design space with described operating constraints. This requirement makes it difficult to apply these methods when feasible parameter ranges or process operating conditions are incomplete or unavailable. In these cases, engineers typically rely on heuristics or domain knowledge to estimate feasible ranges8, 9. These heuristic approaches are highly subjective and may lead to variable optimization results. Mathematical simulation platforms such as IDAES10 and modeling libraries such as Pyomo11 offer flexible environments to build detailed process models, but still depend on users to define variable constraints and smooth objective functions before running optimization. This constraint definition bottleneck significantly limits the application of such optimization techniques, particularly for non-expert users or for novel processes where operational configurations are not well-established.
@@
Recent developments in large language models (LLMs) offer potential approaches to address these limitations. LLMs have demonstrated strong reasoning capabilities in complex problem-solving tasks12, 13 and effective adaptation to previously unseen domains through few-shot learning14. Most remarkably, recent work has shown that LLMs can autonomously conduct scientific research, from hypothesis generation and experimental design to data analysis and manuscript writing15, demonstrating their potential for independent scientific reasoning and discovery.
@@
These general capabilities have translated into significant advances within chemical engineering applications. In synthesis planning, LLMs evaluate synthetic routes based on strategic requirements16, suggesting that domain expertise traditionally requiring years of training could become more accessible through natural language interfaces. For material discovery, LLM-based agents such as ChemCrow17 have successfully planned and executed syntheses using automated laboratory platforms. For control system design18, 19, 20, 21, 22, LLMs enable more flexible process control approaches without the need for specialized programming skills.
@@
Our approach leverages the AutoGen platform23 with OpenAI’s o3 model24 to create a multi-agent system where language model agents collaborate to both generate realistic constraints and iteratively optimize process parameters. The framework is evaluated on a real-world chemical flowsheet - the hydrodealkylation (HDA) process using IDAES10 simulations and assessed across three metrics: cost, yield, and yield-to-cost ratio. We selected IPOPT and grid search as representative benchmarks: IPOPT as a state-of-the-art gradient-based solver widely used in chemical engineering, and grid search as a gradient-free exhaustive method that serves as a baseline for global optimization performance. The results show comparable performance, while requiring fewer iterations and less computational time. This approach could significantly reduce the expertise barrier for process optimization and enable more widespread application of advanced optimization techniques in industrial settings, particularly for emerging processes or retrofitting existing operations where operational constraints are poorly characterized.
@@ 2 Methodology
2.1 Framework Architecture
We propose a multi-agent interactive system using AutoGen’s GroupChat architecture that addresses fundamental limitations of single-agent approaches for complex optimization tasks. Individual agents maintain focused prompts for specific functions, preventing context dilution that occurs when lengthy multi-purpose prompts overwhelm LLM attention mechanisms. Each agent’s specialized role enables direct function calling rather than complex output parsing, reducing implementation complexity and error potential. The modular design allows independent debugging and modification of optimization components, critical for iterative algorithm development and deployment flexibility.
@@ 2.1 Framework Architecture
Figure 1: Architecture of the LLM-based optimization framework illustrating agent workflow. The ContextAgent generates constraints and detailed process overview from basic process info using domain knowledge, while specialized agents within the GroupChat environment handle parameter introduction, validation, optimization suggestions, and IDAES-based simulation evaluation through iterative refinement until optimal solutions are achieved.
@@ 2.2 Agent Roles and Functions
ParameterAgent initiates the process by receiving parameter-value pairs from users as initial starting points. These initial guesses can be arbitrary and may violate feasible bounds, as users typically lack knowledge of the constraint ranges previously generated by the ContextAgent. The ValidationAgent will identify any constraint violations and redirect to the SuggestionAgent for correction.
@@
SimulationAgent executes process evaluation once parameters pass validation. This agent runs a custom objective function that interfaces with a pre-defined python-based IDAES simulation model (detailed IDAES implementation, thermodynamic properties, and kinetic parameters are provided in Supporting Information), calculating performance metrics such as cost, yield, and yield-to-cost ratios for each parameter configuration. Performance results are reported back to the SuggestionAgent, including any simulation convergence issues.
SuggestionAgent serves as the optimization engine, maintaining a comprehensive record of the optimization history. After each iteration, current parameter values and corresponding simulation results are appended to the agent’s memory, providing direct access to both valid bounds and historical trial data. This accumulated knowledge enables the agent to refine and propose future parameter sets based on observed trends and constraint violations, completing the iterative optimization cycle.
@@ 2.3 Optimization Workflow
This structured workflow design addresses three critical challenges in autonomous optimization: (1) constraint enforcement through mandatory validation checkpoints, (2) systematic exploration through memory-guided suggestions, and (3) autonomous termination through trend analysis. Unlike gradient-based methods that require mathematical derivatives or grid search that evaluates fixed discretizations, this agent-based approach leverages reasoning to intelligently navigate parameter spaces, adapting search strategy based on constraint violations and performance feedback. The framework’s efficiency stems from this reasoning-guided approach, enabling convergence in 20-45 iterations compared to thousands required by exhaustive methods, while maintaining solution quality through systematic constraint enforcement and trend analysis.
@@ 2.4 Evaluation Metrics
Annual operating cost serves as the primary economic indicator, minimizing energy consumption through optimal utility management including heating costs and cooling costs for operating units:
[…]
@@
[…]
@@
[…]
These metrics were selected to cover industrial optimization priorities: economic performance (cost), production effectiveness (yield), and overall process efficiency (yield-to-cost ratio), allowing direct comparison with conventional optimization approaches across multiple performance dimensions.
@@ 3.1 Case Study Validation
Traditional optimization methods like IPOPT or grid search cannot be directly applied without explicit constraint sets. This limitation demonstrates the practical need for our ContextAgent, which infers variable bounds using embedded process knowledge. Given this constraint generation capability, we formulate the optimization problem as follows:
$\text{optimize }\phi(x)\quad\text{subject to}\quad x_{i}^{L}\leq x_{i}\leq x_{i}^{U},\quad\forall i=1,\dots,4$ (4)
@@ 3.2 Constraint Generation Quality
To assess the reliability and consistency of autonomous constraint generation, the ContextAgent’s constraint generation capability was evaluated across five independent trials using identical prompts to ensure representative constraint generation and minimize stochastic variance. Table 1 presents the individual trial results and average constraint ranges inferred for the operating conditions across key unit operations H101, F101, and F102. The average ranges shown in the rightmost column were used for subsequent optimization performance comparisons.
Table 1: ContextAgent Constraint Generation Results and Average Ranges Across Five Trials
Unit Variable Trial 1 Trial 2 Trial 3 Trial 4 Trial 5 Average Range
H101 Temp (K) [870, 950] [873, 973] [820, 980] [800, 950] [773, 1023] [827.2, 975.2]
F101 Temp (K) [310, 330] [300, 340] [290, 390] [330, 420] [298, 368] [305.6, 369.6]
F102 Temp (K) [300, 340] [330, 380] [280, 340] [300, 360] [323, 413] [306.6, 366.6]
F102 $\Delta P$ (Pa) [-200000, -50000] [-200000, -50000] [-240000, -10000] [-200000, -10000] [-220000, -20000] [-216000, -28000]
@@
Importantly, despite this observable variation across trials, the ContextAgent consistently generated realistic and engineering-appropriate bounds. For the critical H101 reactor temperature, the average generated range of [827.2, 975.2] K shows a good alignment with industrial practices reported at [773, 977] K25, 8. Although specific industrial ranges for the flash separator temperatures and pressure drop are not readily available in the literature, the close correspondence between agent-generated reactor constraints and established industrial practice demonstrates the agent’s ability to extract field-consistent operational conditions using only basic process information and embedded domain knowledge.
@@ 3.3 Performance Comparison
Optimization Results: To assess the effectiveness of our approach, we compared the LLM framework’s optimization performance against IPOPT (detailed algorithm description is provided in Supporting Information) and grid search benchmarks using the averaged ContextAgent-generated constraints. Grid search was implemented with 10 discretization points per variable, resulting in 10,000 total parameter combinations (10×10×10×10) evaluated across the four decision variables. Table 2 summarizes the optimization results across the three evaluation metrics as defined earlier.
@@
Metric Method H101 Temp (K) F101 Temp (K) F102 Temp (K) $\Delta P$ F102 (Pa) Value
Cost (\$/yr)a LLM 830.0 325.0 360.0 -180000 $5.792\times 10^{5}$
IPOPT 827.2 305.6 306.6 -28314 $5.660\times 10^{5}$
Grid Search 827.2 305.6 306.6 -48444 $5.660\times 10^{5}$
Yield (mol/yr)b LLM 975.0 369.5 366.0 -180000 $1.644\times 10^{7}$
IPOPT 857.5 369.6 344.4 -130340 $1.565\times 10^{7}$
Grid Search 958.8 369.6 339.9 -89333 $1.750\times 10^{7}$
Yield/Cost (mol/\$)b LLM 827.2 369.0 306.7 -28000 17.23
IPOPT 827.2 369.6 306.6 -121039 17.62
Grid Search 827.2 369.6 306.6 -68889 17.62
@@
Performance Analysis: The results shown above demonstrate competitive performance across all three metrics. Grid search consistently achieves the best solutions, benefiting from exhaustive evaluation of the discretized parameter space. However, the LLM-based optimization framework demonstrates far better optimization performance compared to conventional methods. Table 3 shows the number of iterations required for each optimization approach across different metrics.
@@
(a) Convergence Speed (Iterations to Converge)
@@
All computational benchmarks were performed on a MacBook Pro equipped with an Apple M1 chip (8-core, 16 GB RAM).
@@
“Keep the reactor charge just inside the allowable window – the nearer-to-minimum H101 temperature ($\approx 830K$) cuts furnace duty and fuel. Push both flash drums to the warm end of their ranges; warmer condensers mean less refrigeration/brine duty. Relax the second-flash pressure drop toward the mechanical minimum (–30 kPa). Throttling is essentially "free," so choosing the smallest $\Delta P$ avoids excessive vapour generation that would otherwise have to be recompressed or condensed downstream.”
This reasoning demonstrates three key capabilities: correct identification of heating-cooling utility trade-offs, understanding of operating condition impacts on energy consumption, and application of domain-informed heuristics for parameter selection. The agent’s ability to articulate process engineering principles while making parameter adjustments distinguishes this approach from conventional numerical optimizers. While this local cost-minimization strategy may not directly lead to the global optimum, it provides efficient search directions that can be refined through subsequent iterations.
Table 4 demonstrates how the agent’s reasoning translates into systematic parameter adjustments during cost minimization. The optimization process begins with arbitrary user-provided initial guesses that may violate the autonomously generated constraints. As shown, the initial H101 temperature of 600 K violates the average generated range [827.2, 975.2] K, resulting in "Invalid" status until the ValidationAgent-SuggestionAgent loop produces the first feasible solution at iteration 1. The progression reveals the agent’s ability to systematically test different operating strategies and adapt based on performance feedback. Early parameter exploration shows varying cost performance as the agent tests different operating strategies, while subsequent iterations show the framework’s convergence toward more cost-effective operating conditions. By iteration 10, the agent identifies an optimal reactor temperature of 830 K that substantially reduces operating costs, with later iterations focusing on fine-tuning secondary variables while maintaining this temperature setting.
@@
Methodological Advantages: The LLM approach does not require objective functions to be differentiable or continuous, potentially making it applicable to optimization problems with discontinuous or non-differentiable objectives. The scalability advantages become particularly evident when considering problem dimensionality. Grid search suffers from exponentially increasing computational demands as the number of decision variables grows, with our four-variable implementation already requiring over 10 hours wall time despite using only 10 discretization points per variable. While the LLM framework’s scalability with problem complexity remains to be validated, its current efficiency suggests potential advantages for complex optimization scenarios where intelligently guided parameter exploration is essential.

### v2

@@ Abstract
Chemical process optimization is crucial to maximize production efficiency and economic performance. Optimization algorithms, including gradient-based solvers, numerical methods, and parameter grid searches, become impractical when operating constraints are ill-defined or unavailable, requiring engineers to rely on subjective heuristics to estimate feasible parameter ranges. To address this constraint definition bottleneck, we present a multi-agent framework of large language model (LLM) agents that autonomously infer operating constraints from minimal process descriptions, then collaboratively guide optimization using the inferred constraints. Our AutoGen-based agentic framework employs OpenAI’s o3 model, with specialized agents for constraint generation, parameter validation, simulation execution, and optimization guidance. Through two phases: (i) autonomous constraint generation using embedded domain knowledge, and (ii) iterative multi-agent optimization, the framework eliminates the need for predefined operational bounds. Validated on the hydrodealkylation process across cost, yield, and yield-to-cost ratio metrics, the framework demonstrated competitive performance with conventional optimization methods while achieving a 31-fold reduction in wall-time relative to grid search, converging in under 20 minutes and requiring far fewer iterations to converge. Beyond computational efficiency, the framework’s reasoning-guided search demonstrates sophisticated process understanding, correctly identifying utility trade-offs, and applying domain-informed heuristics. Unlike conventional optimization methods like Bayesian Optimization that require predefined constraints, our approach uniquely combines autonomous constraint generation with interpretable, reasoning-guided parameter exploration. Reproducibility analysis across five independent trials demonstrates consistent convergence behavior, while model comparison reveals that reasoning-capable LLM architectures (o3, o1) are essential for successful optimization, with standard models failing to converge effectively. This approach shows significant potential for optimization scenarios where operational constraints are poorly characterized or unavailable, particularly for emerging processes and retrofit applications.
Keywords: large language model, process optimization, process design, agentic framework
@@ 1 Introduction
Efficient optimization of chemical processes is a central objective in process systems engineering, with direct implications for production efficiency and economic performance[1]. Various optimization approaches have been developed to address different aspects of process optimization challenges. Bayesian optimization (BO)[2, 3] is one approach to black-box optimization, though it is computationally expensive and requires substantial function evaluations to converge. Gradient-based solvers remain the dominant methods in process optimization, including Interior Point OPTimizer (IPOPT)[4] and Sparse Nonlinear OPTimizer (SNOPT)[5], which depend on the smoothness and differentiability of the underlying objective function[6]. When these requirements cannot be met, alternative methods such as evolutionary algorithms[7], and derivative-free methods[8] provide viable solutions. Furthermore, exhaustive search techniques[9] such as grid search systematically evaluate all combinations of discretized parameter values throughout the design space.
@@
However, all of these conventional optimization methods, including BO, require a complete and well-defined design space with described operating constraints. This requirement makes it difficult to apply these methods when feasible parameter ranges or process operating conditions are incomplete or unavailable. In these cases, engineers typically rely on heuristics or domain knowledge to estimate feasible ranges[10, 11]. These heuristic approaches are highly subjective and may lead to variable optimization results. Mathematical simulation platforms such as IDAES[12] and modeling libraries such as Pyomo[13] offer flexible environments to build detailed process models and utilize IPOPT as its solver for scalability, speed. These platforms, however, still require users to define appropriate variable bounds and formulate smooth, well-posed objective functions before optimization. This constraint definition bottleneck significantly limits the application of optimization techniques across the board, particularly for non-expert users or for novel processes where operational configurations are not well-established.
@@
Recent developments in large language models (LLMs) offer potential approaches to address these limitations. LLMs have demonstrated strong reasoning capabilities in complex problem-solving tasks[14, 15, 16] and effective adaptation to previously unseen domains through few-shot learning[17]. Most remarkably, recent work has shown that LLMs can autonomously conduct scientific research, from hypothesis generation and experimental design to data analysis and manuscript writing[18], demonstrating their potential for independent scientific reasoning and discovery.
@@
These general capabilities have translated into significant advances within chemical engineering applications[19, 20, 21]. In synthesis planning, LLMs evaluate synthetic routes based on strategic requirements[22], suggesting that domain expertise traditionally requiring years of training could become more accessible through natural language interfaces. For material discovery, LLM-based agents such as ChemCrow[23] have successfully planned and executed syntheses using automated laboratory platforms. For control system design[24, 25, 26, 27, 28], LLMs enable more flexible process control approaches without the need for specialized programming skills.
@@
Our approach leverages the AutoGen platform[29] with OpenAI’s o3 model[30] to create a multi-agent system where language model agents collaborate to both generate realistic constraints and iteratively optimize process parameters. The framework is evaluated on a real-world chemical flowsheet - the hydrodealkylation (HDA) process, using IDAES[12] simulations and assessed across three metrics: cost, yield, and yield-to-cost ratio. We selected IPOPT and grid search as representative benchmarks: IPOPT as a state-of-the-art gradient-based solver widely used in chemical process modeling, and grid search as a gradient-free exhaustive method that provides a global optimization baseline across the full parameter space. The results show comparable performance, while requiring fewer iterations and less computational time. This approach could significantly reduce the expertise barrier for process optimization and enable more widespread application of advanced optimization techniques in industrial settings, particularly for emerging processes or retrofitting existing operations where operational constraints are poorly characterized.
@@ 2 Methodology
2.1 Framework Design
We propose a multi-agent interactive system using AutoGen’s GroupChat architecture[29] that addresses fundamental limitations of single-agent approaches for complex optimization tasks. Individual agents maintain focused prompts for specific functions, preventing context dilution that occurs when lengthy multi-purpose prompts overwhelm LLM attention mechanisms. Each agent’s specialized role enables direct function calling rather than complex output parsing, reducing implementation complexity and error potential. The modular design allows independent debugging and modification of optimization components, critical for iterative algorithm development and deployment flexibility.
@@ 2.1 Framework Design
Figure 1: Multi-agent workflow of the LLM-based optimization framework. The ContextAgent generates constraints and detailed process overview from basic process info using domain knowledge, while specialized agents within the GroupChat environment handle parameter introduction, validation, optimization suggestions, and IDAES-based simulation evaluation through iterative refinement until optimal solutions are achieved.
@@
Here, we document that all computational experiments in this study were performed on a MacBook Pro equipped with an Apple M1 chip (8-core, 16 GB RAM). The framework requires no model training or GPU resources, relying on API-based access to pre-trained language models.
@@ 2.2 Agent Roles and Functions
ParameterAgent initiates the process by receiving parameter-value pairs from users as initial starting points. These initial guesses can be arbitrary and may violate feasible bounds, as users typically lack knowledge of the constraint ranges previously generated by the ContextAgent.
@@
SimulationAgent executes process evaluation once parameters pass validation. This agent runs a custom objective function that interfaces with a pre-defined python-based IDAES[12] simulation model (detailed IDAES implementation, thermodynamic properties, and kinetic parameters are provided in Supporting Information), calculating performance metrics such as cost, yield, and yield-to-cost ratios for each parameter configuration. Performance results are reported back to the SuggestionAgent, including any simulation convergence issues.
SuggestionAgent serves as the optimization engine, maintaining a comprehensive record of the optimization history. After each iteration, current parameter values and corresponding simulation results are appended to the agent’s memory, providing direct access to both valid bounds and historical trial data. This accumulated knowledge enables the agent to refine and propose future parameter sets based on observed trends and constraint violations, completing the iterative optimization cycle. Figure 2 illustrates the SuggestionAgent workflow with simplified representations of the prompt instructions, process overview, and evaluation history, with the complete prompt provided in Supporting Information.
Figure 2: Visual illustration of the workflow of the SuggestionAgent. The agent receives three key inputs: (1) a structured prompt with optimization instructions, (2) process overview including valid constraint ranges, and (3) evaluation history of previous attempts. The SuggestionAgent analyzes these inputs to generate feasible parameter suggestions that satisfy all constraints, which are then validated and refined through iterative multi-agent interactions.
@@ 2.3 Optimization Workflow
This structured workflow design addresses three critical challenges in autonomous optimization: (1) constraint enforcement through mandatory validation checkpoints, (2) systematic exploration through memory-guided suggestions, and (3) autonomous termination through trend analysis. Unlike gradient-based methods that require mathematical derivatives or grid search that evaluates fixed discretizations, this agent-based approach leverages reasoning to intelligently navigate parameter spaces, adapting search strategy based on constraint violations and performance feedback.
@@ 2.4 Evaluation Metrics
Annual operating cost[31] serves as the primary economic indicator, minimizing energy consumption through optimal utility management including heating costs and cooling costs for operating units:
$\rm{Operating\ Cost}=\rm{Heating\ Cost}+\rm{Cooling\ Cost}\quad[\mathdollar/\rm{yr}]$ (1)
@@
$\rm{Yield}=\rm{F_{benzene,reactor}}\times 3600\times 24\times 365\quad[mol/yr]$ (2)
where $F_{benzene,reactor}$ is explicitly defined as the molar flow rate of benzene at the reactor outlet (R101) in units of $mol/s$.
@@
$\mbox{Yield--to--Cost Ratio}=\frac{\rm{Annual\ Yield}}{\rm{Operating\ Cost}}\quad[\rm{mol/\mathdollar}$] (3)
These metrics were selected to cover industrial optimization priorities: economic performance (cost), production effectiveness (yield), and overall process efficiency (yield-to-cost ratio), allowing direct comparison with conventional optimization approaches across multiple performance dimensions. It is important to note that "Operating Cost" is defined as the sum of variable utility costs (Heating Cost + Cooling Cost) and thus excludes fixed charges or capital costs.
@@ 3.1 Case Study Validation
Mathematical optimization methods like IPOPT or grid search cannot be directly applied without explicit constraint sets. This limitation demonstrates the practical need for our ContextAgent, which infers variable bounds using embedded process knowledge. Given this constraint generation capability, we formulate the optimization problem as follows:
$\rm{optimize\ }\phi(x)\quad\rm{subject\ to}\quad x_{i}^{L}\leq x_{i}\leq x_{i}^{U},\quad\forall i=1,\dots,4$ (4)
@@
While BO achieves sample efficiency when optimizing well-behaved continuous functions with well-defined constraints[2] [3], our problem presents characteristics that align poorly with BO’s assumptions. First, constraint generation is part of the optimization problem. Unlike hyperparameter tuning where valid ranges are known, chemical process optimization requires determining feasible operating boundaries based on thermodynamic and kinetic principles as well as equipment limitations. Second, engineers need to understand why specific operating conditions are recommended for safety validation and regulatory compliance. LLMs provide natural language explanations grounded in chemical engineering principles, allowing them to justify parameter suggestions (e.g., ’Increase temperature to 350 °C to improve reaction kinetics while staying below thermal decomposition limits’) and identify violations of physical laws or impractical operating regimes (e.g., conditions causing unintended phase changes or exceeding material temperature limits). In contrast, BO’s parameter selection lacks interpretable explanations and cannot leverage encoded domain knowledge to validate recommendations.
@@ 3.2 Constraint Generation Quality
To assess the reliability and consistency of autonomous constraint generation, the ContextAgent’s constraint generation capability was evaluated across five independent trials using identical prompts to ensure representative constraint generation and minimize stochastic variance. The coefficient of variation (CV) was calculated to quantify the consistency of constraint generation across trials. For each variable $i$, the CV was computed as the average of the lower-bound and upper-bound CVs:
$CV_{i}=\frac{1}{2}\left(\frac{\sigma_{L,i}}{\mu_{L,i}}+\frac{\sigma_{U,i}}{\mu_{U,i}}\right)\times 100\%$
where $\sigma_{L,i}$ and $\mu_{L,i}$ are the standard deviation and mean of the lower bounds, and $\sigma_{U,i}$ and $\mu_{U,i}$ are the standard deviation and mean of the upper bounds across the five trials. Table 1 presents the individual trial results, average constraint ranges, and CV values for the operating conditions across key unit operations H101, F101, and F102. The average ranges were used for subsequent optimization performance comparisons.
Table 1: ContextAgent Constraint Generation Results and Average Constraint Range (with Variation) Across Five Independent Trials
Unit Variable Trial 1 Trial 2 Trial 3 Trial 4 Trial 5 Average Range (CV (%) $\downarrow$ )
H101 Temp (K) [870, 950] [873, 973] [820, 980] [800, 950] [773, 1023] [827.2, 975.2] (4.18)
F101 Temp (K) [310, 330] [300, 340] [290, 390] [330, 420] [298, 368] [305.6, 369.6] (7.49)
F102 Temp (K) [300, 340] [330, 380] [280, 340] [300, 360] [323, 413] [306.6, 366.6] (7.48)
F102 $\Delta P$ (Pa) [-200000, -50000] [-200000, -50000] [-240000, -10000] [-200000, -10000] [-220000, -20000] [-212000, -28000] (40.82)
@@
Quantitatively, the coefficient of variation (CV) analysis showed low variability for temperature-related constraints, below $10\%$, but substantially higher variability for the pressure drop constraint, around $40\%$. This indicates that the temperature bounds were generated more consistently across trials, whereas the pressure constraint was more sensitive to prompt stochasticity and less tightly grounded in the provided textual context.
Importantly, despite this observable variation across trials, the ContextAgent consistently generated realistic and engineering-appropriate bounds. For the critical H101 reactor temperature, the average generated range of [827.2, 975.2] K shows a good alignment with industrial practices reported at [773, 977] K[32, 10]. Although specific industrial ranges for the flash separator temperatures and pressure drop are not readily available in the literature, the close correspondence between agent-generated reactor constraints and established industrial practice demonstrates the agent’s ability to extract field-consistent operational conditions using only basic process information and embedded domain knowledge.
@@ 3.3 Performance Comparison
Optimization Results: To assess the effectiveness of our approach, we compared the LLM framework’s optimization performance against IPOPT (detailed algorithm description is provided in Supporting Information) and grid search benchmarks using the averaged ContextAgent-generated constraints. Grid search was implemented with 10 discretization points per variable, resulting in 10,000 total parameter combinations (10×10×10×10) evaluated across the four decision variables. The achievement metric was calculated to quantify each method’s optimization performance relative to the grid search baseline. For each method, the achievement was computed as the ratio of the objective value obtained by the evaluated method to that obtained by the grid search benchmark:
$\mathrm{Achievement}=\frac{V_{\mathrm{method}}}{V_{\mathrm{grid}}}\times 100\%$
where $V_{\mathrm{method}}$ and $V_{\mathrm{grid}}$ are the objective values achieved by the evaluated optimization method and the grid search baseline, respectively. Table 2 summarizes the optimization results across the three evaluation metrics as defined earlier.
@@
Metric Method H101 Temp (K) F101 Temp (K) F102 Temp (K) F102 $\Delta P$ (Pa) Value Achievement (%) ($\uparrow$)
Cost (\$/yr)a LLM 830.0 325.0 360.0 -180000 $5.792\times 10^{5}$ 97.72
[…]
Yield (mol/yr)b LLM 975.0 369.5 366.0 -180000 $1.644\times 10^{7}$ 93.94
IPOPT 857.5 369.6 344.4 -130340 $1.565\times 10^{7}$ 89.43
[…]
Yield/Cost ($\mathrm{mol/\mathdollar})^{b}$ LLM 827.2 369.0 306.7 -28000 17.23 97.79
IPOPT 827.2 369.6 306.6 -121039 17.62 100.00
Grid Search 827.2 369.6 306.6 -68889 17.62 100.00
@@
Performance Analysis: The results shown above demonstrate competitive performance across all three metrics. Grid search consistently achieves the best solutions, benefiting from exhaustive evaluation of the discretized parameter space. However, the LLM-based optimization framework demonstrates comparable or better performance while being dramatically faster than conventional methods. Table 3 shows the number of iterations required for each optimization approach across different metrics.
@@
(a) Convergence Speed (Iterations)
@@
To assess convergence behavior and reproducibility, we conducted five independent cost minimization runs using identical constraints and process descriptions. The framework demonstrated rapid convergence across runs, indicating consistent optimization behavior. Detailed convergence analysis are provided in Supporting Information.
@@
"Keep the reactor charge just inside the allowable window – the nearer-to-minimum H101 temperature ($\approx 830K$) cuts furnace duty and fuel. Push both flash drums to the warm end of their ranges; warmer condensers mean less refrigeration/brine duty. Relax the second-flash pressure drop toward the mechanical minimum (–30 kPa). Throttling is essentially "free," so choosing the smallest $\Delta P$ avoids excessive vapour generation that would otherwise have to be recompressed or condensed downstream."
This reasoning demonstrates three key capabilities: correct identification of heating-cooling utility trade-offs, qualitative understanding of operating condition impacts on energy consumption, and application of domain-informed heuristics for parameter selection. The agent’s ability to articulate process engineering principles while making parameter adjustments distinguishes this approach from conventional numerical optimizers. While this local cost-minimization strategy may not directly lead to the global optimum, it provides efficient search directions that can be refined through subsequent iterations.
Table 4 demonstrates how the agent’s reasoning translates into systematic parameter adjustments during cost minimization. The optimization process begins with arbitrary user-provided initial guesses that may violate the autonomously generated constraints. As shown, the initial H101 temperature of 600 K violates the average generated range [827.2, 975.2] K, resulting in "Invalid" status until the ValidationAgent-SuggestionAgent loop produces the first feasible solution at iteration 1. The progression reveals the agent’s ability to systematically test different operating strategies and adapt based on performance feedback. Early parameter exploration shows varying cost performance as the agent tests different operating strategies, while subsequent iterations show the framework’s convergence toward more cost-effective operating conditions. By iteration 10, the agent identifies an optimal reactor temperature of 830 K that substantially reduces operating costs, with later iterations focusing on fine-tuning secondary variables while maintaining this temperature setting. The framework’s efficiency stems from this reasoning-guided approach, enabling convergence in 20-45 iterations compared to thousands required by exhaustive methods, while maintaining solution quality through systematic constraint enforcement and trend analysis.
@@
Framework Robustness Across Models: The reasoning capabilities demonstrated above are not universal across all LLMs. To assess the framework’s robustness and reproducibility, we evaluated the cost minimization task using four different models: OpenAI’s o3 and o1 (reasoning models), and GPT-4o and GPT-4.1 (standard models). All models received identical prompts, constraints, and access to optimization history, isolating model reasoning capability as the sole variable.
Figure 4 reveals a significant performance gap between specialized reasoning models and general models. Only o3 and o1 successfully converged to solutions comparable with the grid-search benchmark, with o3 demonstrating superior efficiency by reaching the benchmark solution in 11 iterations compared to o1’s 14 iterations. In contrast, GPT-4o and GPT-4.1 failed to converge, stalling at suboptimal solutions after merely 4 and 5 iterations respectively, despite having access to the same information and agent tools.
Figure 4: Convergence comparison across different LLM architectures for cost minimization. Solid dots indicate where models terminated, with dashed lines extending to show their final achieved cost for comparison. Red dashed line represents grid search’s optimal cost solution of $5.660\times 10^{5}/yr$. Reasoning models (o3 and o1) successfully converge, with o3 requiring 11 iterations and o1 requiring 14 iterations. Standard models (GPT-4o and GPT-4.1) terminate prematurely after 4 to 5 iterations without effective learning.
This performance gap highlights a critical requirement for autonomous optimization: successful operation demands more than access to domain knowledge or the ability to follow instructions. It requires sustained reasoning capabilities that enable multi-step planning, constraint satisfaction, and iterative refinement based on feedback. The superior performance of o3 over o1, despite both being reasoning models, suggests that advances in reasoning depth and consistency directly translate to optimization efficiency. This finding has important implications for framework deployment: while the approach demonstrates feasibility for autonomous constraint generation and optimization, its practical application currently requires access to advanced reasoning-capable models. Future work should investigate strategies to enhance robustness across model architectures, potentially through improved prompt engineering, hybrid approaches combining multiple models, or integration with conventional optimization methods for critical decision steps.
Methodological Advantages: The LLM approach does not require objective functions to be differentiable or continuous, potentially making it applicable to optimization problems with discontinuous or non-differentiable objectives. The scalability advantages become particularly evident when considering problem dimensionality. Grid search suffers from exponentially increasing computational demands as the number of decision variables grows (scaling as $n^{d}$ where n is discretization points and d is number of variables), with our four-variable implementation already requiring over 10 hours wall time despite using only 10 discretization points per variable. Our reasoning-guided approach intelligently navigates the parameter space without exhaustive enumeration, suggesting the framework has potential to handle higher-dimensional problems.
Beyond scalability, the framework’s process-agnostic design positions it well for application across diverse chemical processes. The ContextAgent requires only textual process descriptions to generate constraints, while the SimulationAgent interfaces with any IDAES-compatible model. Extending validation to reactive systems and energy systems would strengthen evidence of broad applicability, likely involving prompt customization to incorporate domain-specific heuristics relevant to each process class. This flexibility in prompt customization, supported by the pre-trained LLMs’ embedded domain knowledge, enables extensions across diverse engineering domains and represents valuable opportunities for future work.